OpenStudy (anonymous):

Rahim rode his motorbike for a distance of 30 km from his house to the General Hospital of Muar. If he incresed his average speed by 10 km / h, the time he takes to complete his journey will be reduced by 30 minutes. Calculate his initial average speed. pls help to solve it..

OpenStudy (michele_laino):

the first journey, we can write: \[d = {v_0}t\] where \(d=30\)Km, and \(v_0\) is the average speed, \(t\) is time needed to complete the journey for second time, we can write this: \[d = \left( {{v_0} + 10} \right)\left( {t - 0.5} \right)\] being \(30\) minutes = \(0.5\) hours so we have to solve this system, with respect to \(v_0\): \[\left\{ \begin{gathered} d = {v_0}t \hfill \\ d = \left( {{v_0} + 10} \right)\left( {t - 0.5} \right) \hfill \\ \end{gathered} \right.\] where \(d=10\) Km

OpenStudy (michele_laino):

oops.. \(d=30\) Km

OpenStudy (anonymous):

tq